OpenStudy (anonymous):

What is your solution set for x + y = 0 x - y + 2 = 0

OpenStudy (anonymous):

i think x+y= 0 x-y=-2 x= -2

OpenStudy (mathhelp346):

What hager did was elimination. Do you understand?

hero (hero):

Sadly, that isn't even correct.

hero (hero):

x+y= 0 x-y=-2 _______ 2x = -2 x = -1 y = 1

OpenStudy (anonymous):

\(\Huge\mathsf{*clap\ clap\ clap*}\)

OpenStudy (anonymous):

Thanks guys lol

OpenStudy (anonymous):

oh sorry i was in rush hero is right